# xTrimoPGLM: Unified 100B-Scale Pre-trained Transformer for Deciphering the Language of Protein

### v1

@@ 1.11.1 Dataset
1.11.2 Supervised Fine-tuning.
@@
1.11.3 Reinforcement Self-Training
@@
1.12 Properties Analysis of Generated Sequences using xTrimoPGLM-100B
@@
Statistical properties of the sampled artificial sequences. Mor Specifically, a diverse set of sequences is sampled using a cross product of temperature ($T$in 0.8, 1.0, 1.2, 1.4, 1.6) and nucleus sampling probability ($P$in 0.5, 0.7, 0.9, 1.0) parameters. For each combination of $T$ and $P$, we sample 600 sequences for the comprehensive sequence analysis.
We present the pairwise sequence identity analysis of generated sequences obtained through various combinations of temperature and nucleus sampling factors, as illustrated in Figure S18(a)(b). We observe that higher nucleus sampling factors and temperatures, which flatten the token probability distribution during the generation process, lead to a broader range of sequence diversity. However, it should be noted that the likelihood of selecting the <eos> token also increases under these conditions. Consequently, higher factors may result in shorter sequences, as illustrated in Figure S18(c)(d). Furthermore, our empirical study suggests that the pre-trained model tends to generate repetitive tokens when the temperature drops below 0.8 and the nucleus sampling factor falls under 0.7, which results in abnormal long sequences. Therefore, we recommend a careful calibration of the hyperparameters, specifically the balance between temperature and nucleus sampling factors, to generate protein sequences that conform to the desired specifications.
@@
Trial & Error of Generated Structures with N-gram penalty. We first produced batches of samples with an n-gram penalty (N-gram=3) to reduce the probability of generating repetitive sequences, as shown in Figure S17. The first row depicts sequences with parameter ($T$=1.0, $P$=1.0, N-gram-penalty=3), while the second row removes the n-gram constraints to reduce long loop disorder regions. We find many examples exhibiting low-complexity sequences (e.g., local repeats), where the predicted structures contain long loop disorder regions. We hypothesize that the n-gram penalty potentially impedes the model’s capacity to generate grammatically correct sequences with ease. Once we remove the n-gram penalty, the generated structures tend to be more natural.
Network Construction. We download SCOPe70_2.08 HH-suite database from HH-suite database website. Then we extracted all sequence clusters in scope70_a3m.ffdata file. Each cluster is used as query to SCOPe70 with hhsearch tool. Each sequence in SCOPe70 is represented as a node. Two nodes are linked if one of them can be searched by the other with at least 20 amino acids and 70% hhsearch probability. For generated sequences, we first search UniClust30 v2021_03 database with hhblits. The output profiles are used as input to search SCOPe70 with hhsearch. Edges are also constructed in a similar way.
1.13 Model FLOPs Comparisons
We conduct a comparative analysis of computational resources utilized by different pre-trained protein language models (Table S4). The parameters detailed in this table are meticulously calculated by implementing the models as per the configurations outlined in their respective source papers and accompanying resources, such as code and model checkpoints. When discrepancies arise between a paper’s theoretical account and its practical application, we favor the metrics provided in the paper. From the right-hand side, the total training tokens are computed by multiplying the training steps, global batch size, and sequence length, given that all models listed are sequence language models. The model’s parameters are estimated directly by following the authors’ released implementations and hyperparameters, with the sum of the training parameters calculated while disregarding tied weights and buffers. The total training computed is estimated by first approximating the FLOPs for one forward propagation (1F) of a single training sample. This is then multiplied by three to account for one forward and one backward propagation without activation recomputation (1F1B). The resulting number is then multiplied by the number of samples used during the entire pre-training process, which is equivalent to the total training tokens divided by the sequence length during pre-training. Only matrix multiplication (matmul) operations are considered in the compute statistics, with other operations such as embedding, element-wise addition, softmax, and layer normalization excluded from the FLOP count. The matmuls considered within the attention block include key, query, and value transformations, attention matrix computation, attention over values, and post-attention linear transformation. Hidden size transformations in the feed-forward block, a projection from hidden dimensions into vocabulary dimensions, and a linear transformation in the language model head (if one exists), are also included in the matmul FLOPs. As an example, if hidden states of size (B, L, D) are multiplied by a weight matrix of size (D, 4D), the resulting FLOPs is BLD4D2 (the factor of 2 accounts for multiplication and addition operations). The total training compute for ProtGPT2 is estimated assuming each A100 GPU performs 120 TFLOPs per second. Consequently, 128 A100 GPUs would achieve approximately 5.3e+21 FLOPs over four days of training.
2 xTrimoPGLM-Ab
2.1 xTrimoPGLM-Ab: OAS Fine-tuning for Antibody
@@ 1.13 Model FLOPs Comparisons
We do not directly fine-tune on xTrimoPGLM-100B, mainly due to limitations in computational budgets and considering the inherent lack of diversity in OAS antibody data, most of which are of similar length and have similar framework areas. Hence, we first pre-train xTrimoPGLM-1B model on the general protein dataset SM Section 1.7 This process undertakes 500B tokens. Since antibodies represent a distinct subset of general proteins, then we finetuned the model using the OAS dataset [^1] Observed Antibody Space (OAS) (?) data. Following the paper, we filter OAS data with IMGT schema (?) and therefore get 678m sequences without disorder and incompletion., comprising 1 billion antibody sequences. Considering that the CDRs are the most important parts of an antibody, we randomly mask one or two whole CDRs for 40% of samples with [sMASK]. A further 40% of the samples undergo a random span masking process, while the remaining 20% are subjected to the MLM objective. We exclude the [gMASK] loss from consideration, as it is not required for downstream antibody-related tasks involving long-text generation. When fine-tuning the xTrimoPGLM-Ab-1B model on OAS data, we decrease the maximum learning rate to 2e-5 and make the model consume 500B tokens with 2,048 samples per batch and the 1,024 input length per sample. It takes about 168 hours to use 128 Nvidia A100 80G GPU cards with mixed precision training. We carry out evaluations on two critical undertakings within the realm of drug discovery including assessing the zero-shot naturalness of antibodies and predicting the structural conformation of antibodies.
2.1.1 Zero-shot Naturalness
In protein design and expression, a crucial step involves filtering out proteins with low expression while retaining those with high naturalness. Perplexity (PPL) given by a protein language model can be used to predict the naturalness of proteins (?, ?, ?). For the GLM objective, PPL is calculated by:
@@ 2.1.1 Zero-shot Naturalness
Results. The results are shown in Table S5. Among these, xTrimoPGLM-Ab-1B surpasses other baselines in two datasets. Moreover, we further fine-tune xTrimoPGLM-Ab-1B with the GLM objective with 30 billion tokens to gain xTrimoPGLM-Ab-1B-GLM. Analogously, we fine-tune it with the MLM objective with 100 billion tokens to get xTrimoPGLM-Ab-1B-MLM. Since the consumed tokens (80% tokens) of the GLM objective is 4 times more than that (20% tokens) of the MLM objective in the pre-training stage, xTrimoPGLM-Ab-1B-MLM is fine-tuned with more tokens than xTrimoPGLM-Ab-1B-GLM for a relatively fair comparison. Consequently, xTrimoPGLM-Ab-1B-GLM and xTrimoPGLM-Ab-1B-MLM keep similar results on Dataset 1 with little difference of AUC on pair test, while they benefit from additional training on Dataset 2, as the AUC scores are improved by 0.02 consistently.
@@
2.1.2 xTrimoPGLM-AbFold: Antibody structure prediction
@@
Model Architecture. Our principal hypothesis is that with an adequately proficient encoder, structure prediction models can accommodate complex structures using shallow Evoformer layers and structure modules. Therefore, compared with the current prevailing folding structures, such as ESMFold, AlphaFold2, we introduce the following modifications to xTrimoPGLM-AbFold: 1) We eliminate MSA and template search modules, as they offer minimal benefit for antibody folding in our pre-training and fine-tuning paradigm; 2) Unlike Alphafold2, which employs 48 blocks of Evoformer, and ESMfold, which utilizes 48 layers of folding trunk, we significantly reduce the number of downstream folding blocks from 48 to 1. The architecture of xTrimoPGLM-AbFold is depicted in Figure S14.
Training Settings. For single-chain structure prediction, we convert protein sequences of length $L$ into the format of [human][chain type]sequence<eos>, and feed it into the xTrimoPGLM-Ab-1B model to obtain the hidden representation $\mathbf{M}$ of the last layer. The information corresponding to [human], [chain type] and <eos> are removed from $\mathbf{M}$, where $\mathbf{M}$ $\in\mathbb{R}^{L\times D}$ and $D$ is the size of the hidden dimension of the xTrimoPGLM-Ab-1B model. Then, we extend $\mathbf{M}$ along its $L$ dimension in a pairwise manner to obtain a tensor $\mathbf{Z}\in\mathbb{R}^{L\times L\times 2D}$ (Figure S14). After that, $\mathbf{M}$ and $\mathbf{Z}$ are fed into a single-block Evoformer module for information fusion and then into the structure module for prediction of the angle and position of each residue. For the VH-VL complex, it should be noted that the input is converted into the format of vh_sequence[linker]vl_sequence, where the [linker] is composed of four groups of residue sequences, each of which is composed of four G residues and one S residue, just like GGGGSGGGGSGGGGSGGGGS.
@@
Results. Each experiment is conducted 5 times with different random seeds and reports the averaged results. As demonstrated in Table S6, xTrimoPGLM-AbFold significantly outperforms all other models, notably xTrimoAbFold—an existing state-of-the-art model—in every metric related to antibody structure prediction. The impressive performance of xTrimoPGLM-AbFold implies that a pre-trained antibody model, when fine-tuned with merely a single additional Evoformer (?) block, can emerge as a leading model for antibody structure prediction, even without the presence of MSA and templates.
Table S7 shows the performance of VH-VL complex in different models. AlphaFold-Multimer, which relies heavily on MSA and template information, outperforms most of protein docking algorithms. However, xTrimoPGLM-AbFold, which does not use any MSA or template information, performs comparable with AlphaFold-Multimer, indicating that xTrimoPGLM-Ab-1B has learned sufficient and rich information on antibodies. Crucially, xTrimoPGLM-AbFold achieves a speedup of 6,300$\times$ over the original AlphaFold-Multimer and 103$\times$ over the faster MSA-searching AlphaFold-Multimer (?), owing to the original AlphaFold-Multimer consumes a long time to search MSA (0.8 hours per sample). When we increase the number of Evoformer blocks to 16, xTrimoPGLM-AbFold attains the best performance on all metrics while still maintaining a 2,400$\times$ speedup than the original AlphaFold-Multimer and 40$\times$ speedup than the accelerated AlphaFold-Multimer. It is noteworthy that only a marginal improvement is attained when the number of Evoformer blocks is increased from 1 to 16, which indicates that xTrimoPGLM -Ab has already captured sufficient information for downstream tasks to predict atom positions with precision.
2.1.3 Antibody-specific Generation using xTrimoPGLM-Ab.
@@
  • CDR3 Short Masking (CSM). This strategy involves masking a partial segment of the CDR3 region. We select the length of the masked region based on a uniform distribution within the interval $[3,6]$. Subsequently, a segment of the CDR3 region is randomly replaced with the [sMASK] token. Upon feeding this modified antibody sequence into xTrimoPGLM-Ab-1B, the masked segment of the CDR3 region undergoes a redesign. The comparison between the conformations of the CDR3-redesigned antibodies and the original sequence is depicted in Figure S16(a).
@@
  • CDR3 Whole Masking (CWM). This strategy involves masking the entirety of the CDR3 region with the [sMASK] token, thus necessitating a de novo design approach. Given the increased complexity of this setting, compared to the CSM, the CWM requires more sophisticated computational models. This method provides a comprehensive and integrative methodology to delve deeper into the complexities of antibody functionality, as shown in Figure S16(b).
@@
  • CDR3 Random Mutation (CRM). This strategy adopts a random mutagenesis approach focusing on specific sites within the CDR3 region. It involves randomly selecting 3-6 positions within the CDR3 domain and subsequently introducing random mutations at these sites. This method can be seen as a stochastic baseline that operates at a comparable edit distance. The result is shown in Figure S16(c).
@@
  • CDR3 Random Retrieval (CRR). This strategy comprises the random substitution of the CDR3 region using sequences from other antibodies present in the SARS-CoV-2 wild-type library. The predicted structures are illustrated in Figure S16(d).
@@
Under the aforementioned settings, we generate a set of 6,000 antibodies via xTrimoPGLM-Ab-1B. Six antibodies are randomly selected as depicted in Figure S16. xTrimoPGLM-AbFold is utilized as the structure prediction model. In response to the observation that using CDR3 short masking tends to generate antibodies closely resembling the ground truth with a small edit distance, we implemented a filter to exclude any antibodies with an edit distance of 2 or less. A series of generated sequences and their corresponding edit distances from the ground truth is presented in Table S9. Importantly, it is noteworthy that both the CSM and CWM policies are capable of generating sequences of varying lengths without resorting to mutations or deletions. In contrast, the sequences generated by the two parallel baselines, CRM and CRR, display considerable disorder, regardless of whether there are few mutations or a complete replacement of the entire CDR3 fragment. Our analysis further identifies a relationship between the edit distance and the structure of the generated antibody’s CDR3 region. Specifically, as the edit distance grows, the organization of the CDR3 region tends to degrade, suggesting that even large generative models currently face limitations.

### v2

@@ 5.1 Dataset
5.2 Supervised Fine-tuning.
@@
5.3 Reinforcement Self-Training
@@
6 Statistical Analysis of Generated Sequences
@@
Statistical properties of the sampled artificial sequences. More specifically, a diverse set of sequences is sampled using a cross product of temperature ($T$in 0.8, 1.0, 1.2, 1.4, 1.6) and nucleus sampling probability ($P$in 0.5, 0.7, 0.9, 1.0) parameters. For each combination of $T$ and $P$, we sample 600 sequences for the comprehensive sequence analysis.
We present the pairwise sequence identity analysis of generated sequences obtained through various combinations of temperature and nucleus sampling factors (Supplementary Figure 20(a)(b)). We observe that higher nucleus sampling factors and temperatures, which flatten the token probability distribution during the generation process, lead to a broader range of sequence diversity. However, it should be noted that the likelihood of selecting the <eos> token also increases under these conditions. Consequently, higher factors may result in shorter sequences (Supplementary Figure 20). Furthermore, our empirical study suggests that the pre-trained model tends to generate repetitive tokens when the temperature drops below 0.8 and the nucleus sampling factor falls under 0.7, which results in abnormal long sequences. Therefore, we recommend a careful calibration of the hyperparameters, specifically the balance between temperature and nucleus sampling factors, to generate protein sequences that conform to the desired specifications.
@@
Trial & Error of Generated Structures with N-gram penalty. We first produced batches of samples with an n-gram penalty (N-gram=3) to reduce the probability of generating repetitive sequences (Supplementary Figure 19). The first row depicts sequences with parameter ($T$=1.0, $P$=1.0, N-gram-penalty=3), while the second row removes the n-gram constraints to reduce long loop disorder regions. We find many examples exhibiting low-complexity sequences (e.g., local repeats), where the predicted structures contain long loop disorder regions. We hypothesize that the n-gram penalty potentially impedes the model’s capacity to generate grammatically correct sequences with ease. Once we remove the n-gram penalty, the generated structures tend to be more natural.
7 xTrimoPGLM-Ab: OAS Fine-tuning for Antibody
@@
We do not directly fine-tune on xTrimoPGLM-100B, mainly due to limitations in computational budgets and considering the inherent lack of diversity in OAS antibody data, most of which are of similar length and have similar framework areas. Hence, we first pre-train xTrimoPGLM-1B model on the general protein dataset SM Section Pre-Training Datasets This process undertakes 500B tokens. Since antibodies represent a distinct subset of general proteins, then we finetuned the model using the OAS dataset [^1] Observed Antibody Space (OAS)  (?) data. Following the paper, we filter OAS data with IMGT schema  (?) and therefore get 678m sequences without disorder and incompletion., comprising 1 billion antibody sequences. Considering that the CDRs are the most important parts of an antibody, we randomly mask one or two whole CDRs for 40% of samples with [sMASK]. A further 40% of the samples undergo a random span masking process, while the remaining 20% are subjected to the MLM objective. We exclude the [gMASK] loss from consideration, as it is not required for downstream antibody-related tasks involving long-text generation. When fine-tuning the xTrimoPGLM-Ab-1B model on OAS data, we decrease the maximum learning rate to 2e-5 and make the model consume 500B tokens with 2,048 samples per batch and the 1,024 input length per sample. It takes about 168 hours to use 128 Nvidia A100 80G GPU cards with mixed precision training. We carry out evaluations on two critical undertakings within the realm of drug discovery including assessing the zero-shot naturalness of antibodies and predicting the structural conformation of antibodies.
7.1 Zero-shot Naturalness
In protein design and expression, a crucial step involves filtering out proteins with low expression while retaining those with high naturalness. Perplexity (PPL) given by a protein language model can be used to predict the naturalness of proteins  (?, ?). For the GLM objective, PPL is calculated by:
@@ 7.1 Zero-shot Naturalness
Results. The results are shown in Supplementary Table 9. Among these, xTrimoPGLM-Ab-1B surpasses other baselines in two datasets. Moreover, we further fine-tune xTrimoPGLM-Ab-1B with the GLM objective with 30 billion tokens to gain xTrimoPGLM-Ab-1B-GLM. Analogously, we fine-tune it with the MLM objective with 100 billion tokens to get xTrimoPGLM-Ab-1B-MLM. Since the consumed tokens (80% tokens) of the GLM objective is 4 times more than that (20% tokens) of the MLM objective in the pre-training stage, xTrimoPGLM-Ab-1B-MLM is fine-tuned with more tokens than xTrimoPGLM-Ab-1B-GLM for a relatively fair comparison. Consequently, xTrimoPGLM-Ab-1B-GLM and xTrimoPGLM-Ab-1B-MLM keep similar results on Dataset 1 with little difference of AUC on pair test, while they benefit from additional training on Dataset 2, as the AUC scores are improved by 0.02 consistently.
@@
7.2 xTrimoPGLM-AbFold: Antibody structure prediction
@@
Model Architecture. Our principal hypothesis is that with an adequately proficient encoder, structure prediction models can accommodate complex structures using shallow Evoformer layers and structure modules. Therefore, compared with the current prevailing folding structures, such as ESMFold, AlphaFold2, we introduce the following modifications to xTrimoPGLM-AbFold: 1) We eliminate MSA and template search modules, as they offer minimal benefit for antibody folding in our pre-training and fine-tuning paradigm; 2) Unlike Alphafold2, which employs 48 blocks of Evoformer, and ESMfold, which utilizes 48 layers of folding trunk, we significantly reduce the number of downstream folding blocks from 48 to 1. The architecture of xTrimoPGLM-AbFold is depicted in Supplementary Figure 17.
Training Settings. For single-chain structure prediction, we convert protein sequences of length $L$ into the format of [human][chain type]sequence<eos>, and feed it into the xTrimoPGLM-Ab-1B model to obtain the hidden representation $\mathbf{M}$ of the last layer. The information corresponding to [human], [chain type] and <eos> are removed from $\mathbf{M}$, where $\mathbf{M}$ $\in\mathbb{R}^{L\times D}$ and $D$ is the size of the hidden dimension of the xTrimoPGLM-Ab-1B model. Then, we extend $\mathbf{M}$ along its $L$ dimension in a pairwise manner to obtain a tensor $\mathbf{Z}\in\mathbb{R}^{L\times L\times 2D}$ (Supplementary Figure 17). After that, $\mathbf{M}$ and $\mathbf{Z}$ are fed into a single-block Evoformer module for information fusion and then into the structure module for prediction of the angle and position of each residue. For the VH-VL complex, it should be noted that the input is converted into the format of vh_sequence[linker]vl_sequence, where the [linker] is composed of four groups of residue sequences, each of which is composed of four G residues and one S residue, just like GGGGSGGGGSGGGGSGGGGS.
@@
Results. Each experiment is conducted 5 times with different random seeds and reports the averaged results. As demonstrated in Supplementary Table 10, xTrimoPGLM-AbFold significantly outperforms all other models, notably xTrimoAbFold—an existing state-of-the-art model—in every metric related to antibody structure prediction. The impressive performance of xTrimoPGLM-AbFold implies that a pre-trained antibody model, when fine-tuned with merely a single additional Evoformer  (?) block, can emerge as a leading model for antibody structure prediction, even without the presence of MSA and templates.
Supplementary Table 11 shows the performance of VH-VL complex in different models. AlphaFold-Multimer, which relies heavily on MSA and template information, outperforms most of protein docking algorithms. However, xTrimoPGLM-AbFold, which does not use any MSA or template information, performs comparable with AlphaFold-Multimer, indicating that xTrimoPGLM-Ab-1B has learned sufficient and rich information on antibodies. Crucially, xTrimoPGLM-AbFold achieves a speedup of 6,300$\times$ over the original AlphaFold-Multimer and 103$\times$ over the faster MSA-searching AlphaFold-Multimer  (?), owing to the original AlphaFold-Multimer consumes a long time to search MSA (0.8 hours per sample). When we increase the number of Evoformer blocks to 16, xTrimoPGLM-AbFold attains the best performance on all metrics while still maintaining a 2,400$\times$ speedup than the original AlphaFold-Multimer and 40$\times$ speedup than the accelerated AlphaFold-Multimer. It is noteworthy that only a marginal improvement is attained when the number of Evoformer blocks is increased from 1 to 16, which indicates that xTrimoPGLM -Ab has already captured sufficient information for downstream tasks to predict atom positions with precision.
7.3 Antibody-specific Generation using xTrimoPGLM-Ab.
@@
  • CDR3 Short Masking (CSM). This strategy involves masking a partial segment of the CDR3 region. We select the length of the masked region based on a uniform distribution within the interval $[3,6]$. Subsequently, a segment of the CDR3 region is randomly replaced with the [sMASK] token. Upon feeding this modified antibody sequence into xTrimoPGLM-Ab-1B, the masked segment of the CDR3 region undergoes a redesign. The comparison between the conformations of the CDR3-redesigned antibodies and the original sequence is depicted in Supplementary Figure 18(a).
@@
  • CDR3 Whole Masking (CWM). This strategy involves masking the entirety of the CDR3 region with the [sMASK] token, thus necessitating a de novo design approach. Given the increased complexity of this setting, compared to the CSM, the CWM requires more sophisticated computational models. This method provides a comprehensive and integrative methodology to delve deeper into the complexities of antibody functionality, as shown in Supplementary Figure 18(b).
@@
  • CDR3 Random Mutation (CRM). This strategy adopts a random mutagenesis approach focusing on specific sites within the CDR3 region. It involves randomly selecting 3-6 positions within the CDR3 domain and subsequently introducing random mutations at these sites. This method can be seen as a stochastic baseline that operates at a comparable edit distance. The result is shown in Supplementary Figure 18(c).
@@
  • CDR3 Random Retrieval (CRR). This strategy comprises the random substitution of the CDR3 region using sequences from other antibodies present in the SARS-CoV-2 wild-type library. The predicted structures are illustrated in Supplementary Figure 18(d).
@@
Under the aforementioned settings, we generate a set of 6,000 antibodies via xTrimoPGLM-Ab-1B. Six antibodies are randomly selected as depicted in Supplementary Figure 18. xTrimoPGLM-AbFold is utilized as the structure prediction model. In response to the observation that using CDR3 short masking tends to generate antibodies closely resembling the ground truth with a small edit distance, we implemented a filter to exclude any antibodies with an edit distance of 2 or less. A series of generated sequences and their corresponding edit distances from the ground truth is presented in Supplementary Table 13. Importantly, it is noteworthy that both the CSM and CWM policies are capable of generating sequences of varying lengths without resorting to mutations or deletions. In contrast, the sequences generated by the two parallel baselines, CRM and CRR, display considerable disorder, regardless of whether there are few mutations or a complete replacement of the entire CDR3 fragment. Our analysis further identifies a relationship between the edit distance and the structure of the generated antibody’s CDR3 region. Specifically, as the edit distance grows, the organization of the CDR3 region tends to degrade, suggesting that even large generative models currently face limitations.
8 Model FLOPs Comparisons
We conduct a comparative analysis of computational resources utilized by different pre-trained protein language models (Supplementary Table LABEL:tab:flos_comparison). The parameters detailed in this table are meticulously calculated by implementing the models as per the configurations outlined in their respective source papers and accompanying resources, such as code and model checkpoints. When discrepancies arise between a paper’s theoretical account and its practical application, we favor the metrics provided in the paper. From the right-hand side, the total training tokens are computed by multiplying the training steps, global batch size, and sequence length, given that all models listed are sequence language models. The model’s parameters are estimated directly by following the authors’ released implementations and hyperparameters, with the sum of the training parameters calculated while disregarding tied weights and buffers. The total training computed is estimated by first approximating the FLOPs for one forward propagation (1F) of a single training sample. This is then multiplied by three to account for one forward and one backward propagation without activation recomputation (1F1B). The resulting number is then multiplied by the number of samples used during the entire pre-training process, which is equivalent to the total training tokens divided by the sequence length during pre-training. Only matrix multiplication (matmul) operations are considered in the compute statistics, with other operations such as embedding, element-wise addition, softmax, and layer normalization excluded from the FLOP count. The matmuls considered within the attention block include key, query, and value transformations, attention matrix computation, attention over values, and post-attention linear transformation. Hidden size transformations in the feed-forward block, a projection from hidden dimensions into vocabulary dimensions, and a linear transformation in the language model head (if one exists), are also included in the matmul FLOPs. As an example, if hidden states of size (B, L, D) are multiplied by a weight matrix of size (D, 4D), the resulting FLOPs is BLD4D2 (the factor of 2 accounts for multiplication and addition operations). The total training compute for ProtGPT2 is estimated assuming each A100 GPU performs 120 TFLOPs per second. Consequently, 128 A100 GPUs would achieve approximately 5.3e+21 FLOPs over four days of training.